OpenStudy (anonymous):

The ratio of the areas of two similar triangles is 4:9. Find the ratio of their corresponding medians.

OpenStudy (anonymous):

2:3

OpenStudy (anonymous):

ratio area=\[(ratio,sides)^{2}\]

OpenStudy (anonymous):

areas of two similar triangles are in the ratio of the squares of their corresponding medians area of triangle1 (median of 1)^2 -------------- = ---------------- area of triangle 2 (median of 2)^2 here 4 (median of i)^2 ---- = --------------- 9 (median of 2)^2 or 2^2 (median of i)^2 ----- = --------------- 3^2 (median of 2)^2 => 2 (median of i) ----- = --------------- 3 (median of 2) so the ratio of medians is 2:3

OpenStudy (anonymous):

pls click the GOOD ANSWER button :))